The total height of a building and the flagpole on the roof is 208 feet.The building is 7 times as tall as the flagpole.How tall is the building?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

X+7X=208
8X=208
X=208/8
X=26 FEET FOR THE FLAG POLE
THUS THE BUILDING IS
208-26=182 FEET TALL.

Answer by stanbon(75887)

The total height of a building and the flagpole on the roof is 208 feet.The building is 7 times as tall as the flagpole.How tall is the building?
---------------
Let the flagpole have measue "x".
Then height of the building is "7x".
EQUATION:
x+7x=208
8x=208
x=26 ft.
7x=182 ft. (height of the building)
Cheers,
Stan H.